# promptbox[1][] colframe=black, colback=white, sharp corners, boxrule=1pt, boxsep=8pt, left=0pt, right=0pt, top=0pt, bottom=0pt, #1

### v1

@@ Abstract.
The “privacy paradox” describes the discrepancy between users’ privacy attitudes and their actual behaviors. Mitigating this discrepancy requires solutions that account for both system opaqueness and users’ hesitations in testing different privacy settings due to fears of unintended data exposure. We introduce an empathy-based approach that allows users to experience how privacy behaviors may alter system outcomes in a risk-free sandbox environment from the perspective of artificially generated personas. To generate realistic personas, we introduce a novel pipeline that augments the outputs of large language models using few-shot learning, contextualization, and chain of thoughts. Our empirical studies demonstrated the adequate quality of generated personas and highlighted the changes in privacy-related applications (e.g., online advertising) caused by different personas. Furthermore, users demonstrated cognitive and emotional empathy towards the personas when interacting with our sandbox. We offered design implications for downstream applications in improving user privacy literacy and promoting behavior changes.
privacy paradox, privacy intervention, decision making, empathy
ccs: Security and privacy Human and societal aspects of security and privacyccs: Human-centered computing Human computer interaction (HCI)
@@ 1. Introduction
Privacy paradox (Norberg et al., 2007; Kokolakis, 2017) is a common phenomenon that refers to the discrepancy between the attitude of users and their actual behaviors in managing their privacy. This inconsistency has been observed in various domains, including social networking service (Hallam and Zanella, 2017; Wu, 2019; Obar and Oeldorf-Hirsch, 2020), online shopping (Beresford et al., 2012; Anic et al., 2019; Bandara et al., 2020), mobile app (Pentina et al., 2016; Barth et al., 2019; Rowe, 2020) and Internet-of-Things (Williams et al., 2016). Prior work in behavioral economics and decision research found several cognitive and behavioral biases that lead to the privacy paradox, such as hyperbolic discounting (Gilovich et al., 2002), the immediate gratification (Acquisti, 2004), and the illusion of control (Brandimarte et al., 2009, 2013).
However, bridging the privacy paradox can be challenging for two reasons. From a system perspective, the inherent opaqueness in the system barriers users from making informed decisions about protecting their privacy. The asymmetric information (Acquisti et al., 2015) provided by the system makes it challenging for users to understand what data is collected and how other parties use it (Mazzia et al., 2012; Yao et al., 2017; Li et al., 2017). Consequently, users are unable to make informed decisions to safeguard their personal data while maintaining the desired level of usability and system utility, such as whether to opt out of certain data collection practices, configure the frequency and granularity of data sharing, or the adoption of privacy-enhancing tools. From the user perspective, the fear of exposing personal data (Zrenner et al., 2019; Spiekermann, 2019; Li et al., 2020) while navigating an opaque system can further discourage users from experimenting with different possible privacy configurations to link their available options of privacy choices to their consequences, thereby reinforcing the system’s opaqueness. Once users share their private data, they will no longer have control over how the other party utilizes the information (Pasquale, 2015). Besides, another barrier that prevents users from meeting their privacy goals is their limited experience and lack of privacy literacy. Lay users are prone to perceive fewer privacy threats compared with technicians (Kang et al., 2015). Thus, even if users sometimes know their privacy goals, they still trade off their privacy for convenience, as they believe that their data is well-protected by the system.
@@
To address the aforementioned challenges and limitations, we present an empathy-based method that allows users to experience and observe the correlation between their privacy data and the system outcomes in a real-time and risk-free environment. In this approach, we use personas that come with synthesized personal data based on real-world privacy incidents. Each persona represents a fictional user (Blomkvist, 2002) with a distinctive biography, demographic information, and a large set of synthesized personal data. For example, here is an exemplary biography of a “tangible” persona:
@@
We validate the proposed approach through a prototype (i.e., Privacy Sandbox) and a study involving 15 participants. The results validated the technical feasibility of our approach to generate artificial personas with realistic synthesized personal data. Our findings imply that users can indeed establish empathy with personas when using the Privacy Sandbox and identify links between the persona’s privacy attributes with the observed system outcomes. The results of the study also offer design implications for using the proposed approach to empower users to acquire privacy knowledge and promote behavior change.
@@
  • Introduces an empathy-based approach that allows users to experience the links between privacy behaviors and system outcomes in a risk-free sandbox environment using artificially generated personas.
  • Validates the viability of the proposed approach through a prototype implementation and empirical studies. The study results confirmed the users’ cognitive and emotional empathy toward the generated personas when interacting with the sandbox in the context of target advertisements.
  • Discusses the design implications of adopting this empathy-based privacy persona approach to empower users to acquire privacy knowledge that leads to behavior change.
@@ 2. Related work
2.1. Empirical studies on privacy paradox
The existence of the privacy paradox is a long-lasting debate in privacy studies. Some researchers regard it as a “myth” (Solove, 2021) because the behavior in the privacy paradox studies pertains to the decision-making about risks in very specific contexts. In contrast, their self-reported privacy concerns are much more general. However, many empirical studies (Acquisti and Gross, 2006; Kokolakis, 2017; Chen et al., 2021) have consistently demonstrated the existence of the privacy paradox across various distinct scenarios.
Most empirical studies measure the gap between users’ privacy attitudes and behaviors, commonly using surveys. For example, Madejski et al. (Madejski et al., 2012) used surveys to gauge privacy attitudes, previous privacy settings, and self-reported sharing intentions on Facebook, identifying potential sharing violations by comparing intentions with settings. Colnago et al. (Colnago et al., 2023) also employed within-subjects surveys, revealing mismatches between attitudes/preferences and behaviors.
@@ 2.1. Empirical studies on privacy paradox
Although surveys are able to explore privacy attitudes, they are not reliable when examining irregular or infrequent privacy behavior (Staddon et al., 2013; Kokolakis, 2017). Consequently, many studies combine surveys with experiments for more reliable behavior data collection. For instance, Norberg et al. (Norberg et al., 2007) assessed willingness to disclose information in surveys and later conducted a field study to compare willingness with actual disclosure, finding significant differences. Barth et al. (Barth et al., 2019) measured privacy concerns through surveys and compared results with participants’ actual behavior, represented by the number of intrusive apps downloaded.
@@
While multiple studies have measured the privacy attitude-behavior gap, only a few have proposed ways to address it. Previous research has examined risk awareness (Dienlin and Trepte, 2015), the privacy calculus  (Park, 2013), and digital nudges (Ioannou et al., 2021) as potential solutions. Sutanto et al. (Sutanto et al., 2013) designed a personalized privacy-safe application that retains users ’ information locally on their smartphones while still providing them with personalized products. Mattson et al.  (Mattson et al., 2023) suggested changing negative attitudes in different functional areas to reduce the intention-behavior gap.
@@
Unlike previous methods relying on surveys and/or experiments for on-the-spot decisions, our approach offers a risk-free environment for users to experience online services by using different generated personas’ data. This allows users to reflect on system outcomes and make more informed decisions, potentially mitigating the privacy paradox gap resulting from ad-hoc decisions.
2.2. Approaches to prompt privacy behavior change
@@
Previous theories on privacy behavior (Ajzen, 1991; Bandura and Walters, 1977; LaRose and Rifon, 2007) emphasize the role of attitudes in shaping behavior. Based on them, we distinguish between two approaches in promoting privacy behavior change: the top-down approach, which primarily seeks to foster privacy attitudes and literacy to indirectly influence behavior, and the bottom-up approach, which directly influences behavior through techniques like nudging. We aim to comprehensively assess their effectiveness and improve their impact on privacy behavior.
@@
Top-down approaches (e.g., privacy education) aim to improve privacy behavior by fostering privacy attitudes (AL MUHANDER et al., 2023; Ebert et al., 2021) and enhancing privacy literacy (Franco and Holzer, 2023a). For instance, Desimpelaere et al. (Desimpelaere et al., 2020) observed that privacy literacy training improves children’s understanding and promotes privacy-protective behavior. Sideri et al. (Sideri et al., 2019) found that university-based education enhanced students’ digital knowledge and privacy awareness on social networking sites. Innovative methods like Franco et al. (Franco and Holzer, 2023b) employed technology-enhanced pedagogical scenarios to involve students in active learning by using their own social media traces. Despite increased awareness and knowledge, top-down approaches may not consistently translate into behavior change. Users often struggle to apply experts’ privacy advice (Das et al., 2019) due to its vagueness (Reeder et al., 2017) and lack of alignment with their specific needs and contexts (Zou et al., 2020).
@@
Bottom-up approaches (e.g., digital nudging) alter user privacy behavior by guiding their choices in digital environments using techniques such as visual cues (Schneider et al., 2018), information presentation (Carpenter et al., 2018), default settings (Baek et al., 2014), and incentives (AN et al., 2021). These nudges can reduce data disclosure and influence privacy choices in a short period (Carpenter et al., 2018; Eling et al., 2016). While privacy nudges can alter privacy behavior quickly, they often lead to temporary effects as they may not necessarily improve users’ long-term privacy literacy or decision alignment with their attitudes.
@@ 2.3. The use of empathy in user experience design and persuasive design
In UX field, empathic design (Kouprie and Visser, 2009) aims to enable designers to ‘step into the user’s shoes’ and ‘walk the user’s walk’, thereby crafting products that align with user needs. A solid foundation for comprehending empathy in design research is established through the exploration of philosophy, psychology, and neuroscience literature. Surma-Aho et al.  (Surma-Aho and Hölttä-Otto, 2022) offered a comprehensive review of empathy’s role in design. A pertinent framework, proposed by Hess and Fila (Hess and Fila, 2016), defines empathy along two axes: affective experiences vs. cognitive processes and self-oriented vs. other-oriented perspectives, yielding four dimensions:
@@
Our work draws inspiration from these empathic understanding frameworks and design methods, aiming to investigate whether users can develop empathy toward the generated personas and whether this influences their acquisition of privacy knowledge.
In persuasive design, stimulating empathy is a crucial technique (Caraban et al., 2019). Previous studies have harnessed the malleable nature of empathy to promote prosocial behavior  (Crowley et al., 2011; Sierksma et al., 2015; Taylor et al., 2019). For example, S.H. Taylor et al. (Taylor et al., 2019) found that embedding empathy nudges in social media posts can encourage bystander intervention for cyberbullying victims. Additionally, many researchers have found that designing with empathy can encourage the natural empathetic behavior of people who have existing social ties or shared interests (Taylor et al., 2019; Preece, 1999) A typical example is that VR can enhance cognitive empathy by emphasizing user similarities (Shin, 2018).
@@
While various measurement scales exist to quantify empathy in psychology (Baron-Cohen, 2009; Davis et al., 1980; Hogan, 1969; Escalas and Stern, 2003), our focus is on the relationship between users and personas rather than personal characteristics. We derive inspiration from approaches used to measure designers’ empathy during the design process, including indicators like empathic expressions, personal experiences, respectful questioning, and discussing user facts (Van Rijn et al., 2011). Our empathy measurement approach combines self-report methods and integrates established theoretical frameworks of empathy.
@@ 3.1. Overview
Figure 1. An empathy-based approach where users interact with online services with different personas in a risk-free sandbox without leaking their real personal data. Users can observe and experience the causal effect between their privacy configurations/behaviors and system outcomes, acquire privacy knowledge, and translate the knowledge into actual behavior.
@@
    Emotional Resonance: When users encounter privacy incidents when acting as personas, we believe that users can feel the emotion that the persona would have felt (e.g., frustration or anger from privacy violations; joy from apt personalized recommendations). Users will also realize how they may feel when they encounter a similar incident.
@@
Our approach has two core phases: persona generation and personal data replacement. Persona generation involves constructing personas with sufficient detail. Once completed, in personal data replacement, users will use personas to explore various privacy settings and online services from the perspective of a particular persona. An overview of this methodology can be seen in Fig. 2. In the following sections, we explain the two steps in detail, followed by introducing a prototype that integrates both steps—the Privacy Sandbox.

### v2

# An Empathy-Based Sandbox Approach to Bridge the Privacy Gap among Attitudes, Goals, Knowledge, and Behaviors

@@ Abstract.
Managing privacy to reach privacy goals is challenging, as evidenced by the privacy attitude-behavior gap. Mitigating this discrepancy requires solutions that account for both system opaqueness and users’ hesitations in testing different privacy settings due to fears of unintended data exposure. We introduce an empathy-based approach that allows users to experience how privacy attributes may alter system outcomes in a risk-free sandbox environment from the perspective of artificially generated personas. To generate realistic personas, we introduce a novel pipeline that augments the outputs of large language models (e.g., GPT-4) using few-shot learning, contextualization, and chain of thoughts. Our empirical studies demonstrated the adequate quality of generated personas and highlighted the changes in privacy-related applications (e.g., online advertising) caused by different personas. Furthermore, users demonstrated cognitive and emotional empathy towards the personas when interacting with our sandbox. We offered design implications for downstream applications in improving user privacy literacy.
privacy awareness, privacy intervention, privacy literacy, empathy, sandbox, generated personas
journalyear: 2024copyright: acmlicensedconference: Proceedings of the CHI Conference on Human Factors in Computing Systems; May 11–16, 2024; Honolulu, HI, USAbooktitle: Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24), May 11–16, 2024, Honolulu, HI, USAdoi: 10.1145/3613904.3642363isbn: 979-8-4007-0330-0/24/05ccs: Security and privacy Human and societal aspects of security and privacyccs: Human-centered computing Human computer interaction (HCI)
@@ 1. Introduction
Managing privacy to achieve individuals’ privacy goals is challenging (Acquisti et al., 2020), as evidenced by the discrepancy between people’s privacy attitudes and their actual behaviors (Norberg et al., 2007; Kokolakis, 2017). Such inconsistency has been observed in various domains, such as social networking service (Hallam and Zanella, 2017; Wu, 2019; Obar and Oeldorf-Hirsch, 2020), online shopping (Beresford et al., 2012; Anic et al., 2019; Bandara et al., 2020), mobile app (Pentina et al., 2016; Barth et al., 2019; Rowe, 2020) and Internet-of-Things (Williams et al., 2016).
Bridging the gap among privacy attitudes, goals, knowledge, and behavior can be difficult for two reasons. From a system perspective, the inherent opaqueness in the system prevents users from making informed decisions about protecting their privacy. The asymmetric information (Acquisti et al., 2015) provided by the system makes it difficult for users to understand what data is collected and how other parties use it (Mazzia et al., 2012; Yao et al., 2017; Li et al., 2017). Consequently, users are unable to make informed decisions to safeguard their personal data while maintaining the desired level of usability and system utility, such as whether to opt out of certain data collection practices, configure the frequency and granularity of data sharing, or the adoption of privacy-enhancing tools. From the user perspective, the fear of exposing personal data (Zrenner et al., 2019; Spiekermann, 2019; Li et al., 2020) while navigating an opaque system can further discourage users from experimenting with different possible privacy configurations to link their available options of privacy choices to their consequences, thereby reinforcing the system’s opaqueness. Once users share their private data, they will no longer have control over how the other party utilizes the information (Pasquale, 2015). In addition, another barrier that prevents users from meeting their privacy goals is their limited experience and lack of privacy literacy (Solove, 2021). Lay users are prone to perceive fewer privacy threats compared to technicians (Kang et al., 2015). Thus, even if users sometimes know their privacy goals, they still trade off their privacy for convenience, as they believe that their data are well protected by the system.
@@
Motivated by the aforementioned challenges and limitations of current approaches, we present an empathy-based method that allows users to experience and observe the correlation between their privacy data and the system outcomes in a real-time and risk-free environment. In this approach, we use personas that come with synthesized personal data based on real-world system outcomes. Each persona represents a fictional user (Blomkvist, 2002) with a distinctive biography, demographic information, and a large set of synthesized personal data. For example, here is an exemplary biography of a “tangible” persona:
@@
We validate the proposed approach through a prototype (i.e., Privacy Sandbox) and a study involving 15 participants. The results validated the technical feasibility of our approach to generate artificial personas with realistic synthesized personal data. Our findings imply that users can indeed establish empathy with personas when using the Privacy Sandbox and identify links between the persona’s privacy attributes with the observed system outcomes. The results of the study also offer design implications for using the proposed approach to empower users to acquire privacy knowledge. Contributions of this work represent a first step towards empowering users to empathize with generated artificial personas, understand their own privacy goals, and gain privacy knowledge. Our broader motivation is to use acquired privacy knowledge to encourage changes in privacy behavior, ultimately bridging the gap between attitudes, goals, knowledge, and actual behavior.
@@
  • Introduces an empathy-based approach that allows users to experience the links between privacy attributes and system outcomes in a risk-free sandbox environment using artificially generated personas.
  • Validates the viability of the proposed approach through a proof-of-concept implementation and empirical studies. The study results confirmed the users’ cognitive and emotional empathy toward the generated personas when interacting with the sandbox in the context of target advertisements.
  • Discusses the design implications of adopting this empathy-based privacy persona approach to empower users to acquire privacy knowledge that leads to behavior change in the future.
@@ 2. Related work
2.1. Empirical studies on the privacy attitudes-behavior gap
@@
Most empirical studies measure the gap between users’ privacy attitudes and behaviors, commonly using surveys. For example, Madejski et al. (Madejski et al., 2012) used surveys to gauge privacy attitudes, previous privacy settings, and self-reported sharing intentions on Facebook, identifying potential sharing violations by comparing intentions with settings. Colnago et al. (Colnago et al., 2023) also employed within-subjects surveys, revealing mismatches between attitudes/preferences and behaviors. Although surveys are able to explore privacy attitudes, they are not reliable when examining irregular or infrequent privacy behavior (Staddon et al., 2013; Kokolakis, 2017). Consequently, many studies combine surveys with experiments to collect more reliable behavior data. For instance, Norberg et al. (Norberg et al., 2007) assessed willingness to disclose information in surveys and later conducted a field study to compare the willingness with actual disclosure, finding significant differences. Barth et al. (Barth et al., 2019) measured privacy concerns through surveys and compared the results with the actual behavior of the participants, represented by the number of downloaded intrusive apps.
@@
Although multiple previous studies have measured the privacy attitude-behavior gap, only a few have proposed ways to address it. Previous research has examined risk awareness (Dienlin and Trepte, 2015), the privacy calculus  (Park, 2013), and digital nudges (Ioannou et al., 2021; Story et al., 2022) as potential solutions. Sutanto et al. (Sutanto et al., 2013) designed a personalized privacy-safe application that retains user information locally on their smartphones while still providing them with personalized products. Mattson et al.  (Mattson et al., 2023) suggested changing negative attitudes in different functional areas to reduce the intention-behavior gap.
The inconsistency between privacy attitudes and behaviors has been framed by some researchers as the “privacy paradox” (Norberg et al., 2007). As summarized by Solove (Solove, 2021), research supporting the privacy paradox argues that actual privacy behaviors better indicate individual’s true privacy preferences compared to self-reported privacy attitudes. This privacy paradox results from the distortion of privacy behavior. However, opponents either argue that this is not a privacy issue or that the privacy paradox does not exist. Some studies diverted the phenomenon from the privacy domain and regarded it as a trust issue (Lutz and Strathoff, 2014; AlAbdali et al., 2021). Other researchers refuted the concept of the privacy paradox. For example, Martin (Martin, 2020) found that consumers keep strong privacy expectations even after the disclosure of information. Solove (Solove, 2021) argued that the behaviors in privacy paradox studies pertain to specific contexts, while the stated privacy concerns are much more general. To explain this mixed view, Acquisti et al. (Acquisti et al., 2020) explained that it is due to researchers’ different definitions and interpretations of the “paradox”. Despite the controversy over the term, the evidence supporting the gap between the mental state of privacy and actual behavior is strong (Acquisti and Gross, 2006; Madejski et al., 2012; Adjerid et al., 2018; Barth et al., 2019), which is the main motivation of this work.
@@
Unlike previous methods relying on surveys and/or experiments for on-the-spot decisions, our approach offers a risk-free environment for users to experience online services by using different generated personas’ data. This allows users to reflect on system outcomes and make more informed decisions, potentially mitigating the privacy attitudes-behavior gap resulting from ad-hoc decisions.
2.2. Approaches to enhance privacy awareness and behavior
@@
Previous theories on privacy behavior (Ajzen, 1991; Bandura and Walters, 1977; LaRose and Rifon, 2007) emphasize the role of attitudes in shaping behavior. Based on them, we distinguish two approaches in improving privacy knowledge and promoting privacy behavior change: the top-down approach, which primarily seeks to foster privacy attitudes and literacy to indirectly influence behavior, and the bottom-up approach, which directly influences behavior through techniques like nudging. We aim to comprehensively assess their effectiveness and improve their impact on privacy behavior.
@@
Top-down approaches (e.g., privacy education) aim to improve privacy behavior by fostering privacy awareness (AL MUHANDER et al., 2023; Ebert et al., 2021) and enhancing privacy literacy (Franco and Holzer, 2023a). For instance, Desimpelaere et al. (Desimpelaere et al., 2020) observed that privacy literacy training improves children’s understanding and promotes privacy-protective behavior. Sideri et al. (Sideri et al., 2019) found that university-based education enhanced students’ digital knowledge and privacy awareness on social networking sites. Innovative methods like Franco et al. (Franco and Holzer, 2023b) employed technology-enhanced pedagogical scenarios to involve students in active learning by using their own social media traces. Despite increased awareness and knowledge, top-down approaches may not consistently translate into behavior change. Users often struggle to apply experts’ privacy advice (Das et al., 2019) due to its vagueness (Reeder et al., 2017) and lack of alignment with their specific needs and contexts (Zou et al., 2020).
@@
Bottom-up approaches (e.g., digital nudging) alter user privacy behavior by guiding their choices in digital environments using techniques such as visual cues (Schneider et al., 2018), information presentation (Carpenter et al., 2018), default settings (Baek et al., 2014), and incentives (AN et al., 2021; Barbosa et al., 2019). These nudges can reduce data disclosure and influence privacy choices in a short period (Carpenter et al., 2018; Eling et al., 2016). While privacy nudges can alter privacy behavior quickly, they often lead to temporary effects as they may not necessarily improve users’ long-term privacy literacy or decision alignment with their attitudes.
@@ 2.3. The use of empathy in user experience design and persuasive design
In UX field, empathic design (Kouprie and Visser, 2009) aims to enable designers to “step into the user’s shoes” and “walk the user’s walk”, thereby crafting products that align with user needs. A solid foundation for comprehending empathy in design research is established through the exploration of the literature on philosophy, psychology, and neuroscience. Surma-Aho et al.  (Surma-Aho and Hölttä-Otto, 2022) offered a comprehensive review of empathy’s role in design. A pertinent framework, proposed by Hess and Fila (Hess and Fila, 2016), defines empathy along two axes: affective experiences vs. cognitive processes and self-oriented vs. other-oriented perspectives, yielding four dimensions:
@@
Our work draws inspiration from these frameworks and design methods of empathic understanding, with the aim of investigating whether users can develop empathy toward the generated personas and whether this influences their acquisition of privacy knowledge.
In persuasive design, the stimulation of empathy is a crucial technique (Caraban et al., 2019). Previous studies have used the malleable nature of empathy to promote prosocial behavior  (Crowley et al., 2011; Sierksma et al., 2015; Taylor et al., 2019). For example, Taylor et al. (Taylor et al., 2019) found that embedding empathy nudges in social media posts can encourage bystander intervention for cyberbullying victims. Additionally, many researchers have found that designing with empathy can encourage the natural empathetic behavior of people who have existing social ties or shared interests (Taylor et al., 2019; Preece, 1999) A typical example is that VR can enhance cognitive empathy by emphasizing user similarities (Shin, 2018).
@@
Although there are various measurement scales to quantify empathy in psychology (Baron-Cohen, 2009; Davis et al., 1980; Hogan, 1969; Escalas and Stern, 2003), our focus is on the relationship between users and personas rather than personal characteristics. We derive inspiration from approaches used to measure designers’ empathy during the design process, including indicators such as empathic expressions, personal experiences, respectful questioning, and discussing user facts (Van Rijn et al., 2011). Our empathy measurement approach combines self-report methods and integrates established theoretical frameworks of empathy.
@@ 3.1. Overview
Figure 1. An empathy-based approach where users interact with online services with different personas in a risk-free sandbox without leaking their real personal data. Users can observe and experience the causal effect between their privacy configurations/behaviors and system outcomes, acquire privacy knowledge, and translate the knowledge into actual behavior. Note that this is a general framework. We leave the studies of changes in privacy behaviors as future work.
@@
    Emotional Resonance: When users encounter system outcomes based on the personal data when acting as personas, we believe that users can feel the emotion that the persona would have felt (e.g., frustration or anger from privacy violations; joy from apt personalized recommendations). Users will also realize how they may feel when they encounter a similar incident.
@@
Our approach has two core phases: persona generation and personal data replacement. Persona generation involves constructing personas with sufficient detail. Once completed, in personal data replacement, users will use personas to explore various privacy settings and online services from the perspective of a particular persona. An overview of this methodology can be seen in Fig. 2. Our focus on online advertisements as an example domain of system outcomes is motivated by their broad impact on everyday life, privacy sensitivity, and potential to raise privacy concerns (Zeng et al., 2021). Targeted ads, a major component of online ads, usually track the user’s online behaviors and personal information, but users often lack awareness of the specific privacy data that these ads automatically collect (Ur et al., 2012; Kang et al., 2015). Note that online ads are just one possible manifestation of system outcomes. Other possible system outcomes include social media feeds, personalized news recommendations, and algorithmic decision-making, as detailed in Section 6.6.4. We anticipate that our approach can be applicable to a wider variety of different system outcomes. In the following sections, we explain the two steps in detail, followed by introducing a prototype that integrates both steps—the Privacy Sandbox.